1.     CONS, Inc. common stock is currently selling for $10 per share, is expected to pay a $1.70 dividend, and sell for $14.30 one year from now. What are the dividend yield, growth rate, and total rate of return? Please share details.


1
Expert's answer
2021-04-16T13:07:53-0400

1)  Dividend Yield Growth= (Dividend per share/Market value of stock) * 100%


($1.70/$10) * 100%= 17%



2) Growth Rate= (Targeted future of stock - Present value of stock)/ Present value of stock * 100%


       ($14.30-$10)/$10 * 100%= 43%



3) Total Rate of Return= [(Targeted future value of stock - original stock value) + Dividends]/ Original stock value * 100%


      [($14.30- $10)+$1.70]/$10 * 100%= 60%



The company’s dividends will achieve a growth rate of 43% and will pay out 17% in dividends relative to its stock price. The company’s net gain over the year stood at 60%.